# A Survey of LLM-Driven AI Agent Communication: Protocols, Security Risks, and Defense Countermeasures

### v3

@@ II-C Detailed Comparison with the Most Relevant Surveys
Survey [318] is one of the surveys with the highest relevance score because it focuses on agent communication and analyzes related protocols. However, it still has significant differences from our survey. First, it lacks some critical protocols like AG-UI, ACP-AgentUnion, ACN, Agent Protocol, API Bridge Agent, and Function Calling. Second, it lacks the analysis of security threats. Third, its defenses prospect is limited. Survey [147] focuses on the influences of MCP. As a result, it lacks other protocols, the illustration of agent communication, and security-related discussion. Similarly, the survey [252] also focuses on MCP. It lacks illustrations of other protocols and comprehensive security analyses. Survey [233] comprehensively introduces A2A, lacking discussion about other protocols and security analyses. Survey [67] detailed discussed MCP, A2A, ANP, and ACP(-IBM). It also partially analyzed related security risks and defenses. However, there still lacks other protocols, in-depth security analyses, and systematic illustration of agent communication. Hou et al. [111] discussed the security risks of MCP. They did not consider other protocols and the high-level overview of agent communication.
III LLM-Driven AI Agents
In this section, we review the entire lifetime from LLM to LLM-driven AI agent. Our goal is to help beginners quickly figure out agents, their characteristics, relationships, and applications.
III-A Large Language Model
@@
Large Language Model (LLM) is a new type of artificial intelligence (AI) model trained on large-scale text corpora to understand and generate human language [211]. Once it came out, LLMs have demonstrated unprecedented capabilities across a wide range of domains, including but not limited to natural language understanding and generating [330], logic reasoning [225, 356, 300], code generation [344], and translation [222]. These remarkable performances can be attributed to two major factors. One is that LLMs are built upon a powerful architecture known as the Transformer [276], which effectively models and captures contextual dependencies between tokens and dynamically weighs the importance of different parts of the input. The other key factor, perhaps the most important one, is the massive scales of LLMs that far exceed traditional AI models. When model parameters surpass certain thresholds, LLMs exhibit emergent abilities [299], referring to unexpected capabilities that do not appear in smaller models. As shown in Table II, the parameter scale of an LLM can be hundreds or thousands of times that of traditional AI models.
@@
| FNN | LeNet-5 [157] | 1998 | 60K |
| RNN | Elman Net [69] | 1990 | 100K |
| LSTM | LSTM [109] | 1997 | 1–10M |
| CNN | ResNet-50 [104] | 2015 | 25M |
| CNN | AlexNet [148] | 2012 | 60M |
| CNN | VGG-16 [251] | 2014 | 138M |
| GAN | DCGAN [228] | 2016 | 4M |
| GNN | GCN [142] | 2017 | 23K |
| Autoencoders | DAE [277] | 2008 | 100K |
| Autoencoders | VAE [141] | 2013 | 1M |
| Transformer | GPT-3 [29] | 2020 | 175B |
| Transformer | PaLM [43] | 2022 | 540B |
| Transformer | GPT-4 [5] | 2023 | 1T |
| Transformer | DeepSeek-V3 [176] | 2024 | 671B |
| Transformer | DeepSeek-R1 [94] | 2025 | 671B |
III-B LLM-Driven AI Agents
@@
Perception module. To automatically finish a specified task, agents need the ability to perceive the real-world environment. For example, the autonomous driving agent needs to sense road conditions in real time so as to take actions such as avoiding, driving, or braking [195, 190]. The type of perception ability depends on the domain for which the agent is designed. For instance, an autonomous driving agent needs the ability of visual or radar perception [258, 321], while a code-generating agent may not require such functions [116, 124].
Memory module. The processing of real-world tasks also requires a strong memory. Agents need to have long-term memory to store complex instructions, knowledge, environment interaction history, or other data that may be required in future steps [102, 191, 328]. This usually requires external storage resources to assist the brain, such as databases or memory sharing [80, 82]. In contrast, LLMs do not have such excellent memory ability. Their memory is short-term, which only lasts for rounds of conversations [284, 357].
Reasoning and planning module. LLM acts as the brain of agents due to its excellent capability of reasoning and planning. It intercepts the instructions from users and automatically decomposes the received task into multiple feasible steps [291, 127, 139, 253]. Then, it selects the best plan from different candidates [136, 354, 113]. Besides, it also revises strategies based on environmental feedback, mitigating errors like code bugs or logical inconsistencies [361, 275, 234, 305]. For example, when the autonomous driving module finds that the barrier is closer, it will change the plan to slow down or detour.
Tool module. The tool module is responsible for deeply integrating external resources with the cognitive capabilities of the agent, enabling it to perform complex operations beyond the native capabilities of LLM [331, 304, 167, 186]. For example, through predefined functional interfaces and protocols, a math agent is able to invoke the external computation libraries and symbolic solvers to help it solve mathematical problems [90].
Action module. The action module is the core hub for interaction with the environment. It is responsible for converting the decisions made by LLMs into executable physical or digital operations and obtaining feedback [290, 345]. This module ensures the executability of instructions through structured output control. For example, it immediately stops generating when LLMs generate a complete action description to avoid redundant output interfering with subsequent parsing.
@@
III-C Comparison Between Agents and LLMs
@@
High Autonomy. LLMs can only passively react to the user prompts and then generate responses. They are unable to plan or execute tasks independently. Besides, the response quality highly relies on the prompt skill [302, 85, 203, 178, 68, 33, 360], which seriously affects the user experience. In contrast, agents possess independent capabilities for task decomposition, strategy adjustment, and external tool invocation, which breaks through the passive mode of LLMS and is highly autonomous.
Flexible Multimodal interaction. LLMs have limited capability of handling multimodal inputs, such as text and pictures [240, 338, 342, 360, 155, 363]. Besides, their outputs are also mainly single-modal (e.g., text-only or picture-only), lacking the ability to actively invoke tools to perform physical actions or generate multimodal content. In contrast, agents overcome these drawbacks by deploying multimodal perception frameworks and tool invocation interfaces. They can realize interactions with complex environments, including vision, text, voice, and other physical elements.
Abundant Tool invocation. LLMS usually passively invoke a single tool (such as Function Calling [213]) through predefined API interfaces and can only perform fixed operations as instructed (e.g., calling the weather API to answer queries [337]). In contrast, agents have the capability of active decision-making. They can independently select, combine, and dynamically adjust multiple tools, such as connecting crawlers, databases, and visualization tools, to generate responses [107].
Better Hallucination inhibition. LLMs suffer from a serious problem called hallucination, which refers to that LLMs are likely to generate non-existent knowledge [348, 93, 272, 118, 315, 169]. LLMs mainly rely on the knowledge internalization of training data, making them prone to hallucinations when facing uncovered domains or outdated information. In contrast, agents are able to reduce the error rate by integrating multiple techniques such as Retrieval Augmentation Generation (RAG) [84, 162, 353] or other methods, which can align the action of agents [268, 79].
Dynamic adaptability. Essentially, LLMs are static models whose knowledge is fixed at the training phase. Although techniques such as fine-tuning [334, 175, 115] or model editing [322, 288, 341, 283, 171] reduce the training cost significantly, LLMs still cannot adapt to real-time events well. In contrast, agents are equipped with techniques like online web search, database query, or real-time sensors, which enable them to dynamically adapt to the changes of real-time environments and information.
@@
Worse Security. Agents have WORSE security than LLMs, which is a major weakness of agents. This is because LLMs are only capable of outputting text. Even if the outputs contain illegal or discriminatory content, their influence on the real world is limited. In contrast, since agents are endowed with the ability to invoke tools, they can cause substantial damages to the real world, including but not limited to maliciously/wrongly operating machines, poisoning databases, and paralyzing the system. As a result, it is necessary to concentrate more on the security of agents.
III-D Agent Applications
@@
Scientific Research. Agents are increasingly embedded into the research workflow, enhancing ideation, automation, and discovery. Their contributions span multiple disciplines, such as mathematics [161, 55, 312, 285], chemistry [239, 50, 42, 27], biological sciences [313, 309, 177], and materials Science [151, 219, 197]
@@
Technical and Engineering Systems. Agents play a growing role in engineering domains, improving automation, systems, and software intelligence. For example, agents are widely used in software engineering, assisting in code generation, bug localization, verification, and system configuration [280, 37, 125, 114, 138, 187]. Besides, agents are also popular in game development and simulation [201, 255]. Embodied intelligence is also another hot topic [36, 194].
@@
Social Governance and Public Services Agents are increasingly deployed in sectors focused on public service and human welfare. For example, agents are now widely used in the legal field to help draft contracts, review legal documents, check compliance rules, and analyze cases [270, 121, 199, 269]. Besides, other fields, such as financial services [164, 319, 212, 329, 98, 97, 73], education [287, 259, 193, 59, 64, 206], and healthcare [35, 298, 74, 238, 23, 123, 289, 159, 346], are also actively integrating agents into their respective practices.
@@
Overall, it can be seen that agents are being widely applied in all walks of life, greatly promoting the development of productivity. More importantly, the application of agents is still in its infancy and has an even greater space for development in the future. It is estimated that the agent market will grow at a rate of 40% annually and is expected to exceed 216.8 billion dollars by 2035 [15].
III-E Takeaways
@@
Agents show multiple advantages over LLMs on multiple metrics, such as richer perception ability, stronger learning ability, and higher adaptivity. Now, to improve the service quality, agents are evolving towards refinement to obtain professional skill in a small domain, no longer pursuing the comprehensive capabilities like LLM. LLMs are more like an intermediate transitional form of the future intelligence, while agents are the next stage of development direction of artificial intelligence. It can be foreseen that they will ultimately become indispensable components of future production ecosystems and daily life. However, agents show worse security than LLMs due to their capability of executing tools. As a result, studying the security of agent communication is significant to the AI ecosystem.
IV Agent Communication Overview
IV-A Motivation: The Demand for Agent Communication
Although the advantages of agents in various fields have become increasingly obvious, their development has also encountered new obstacles, which have given rise to the demand for agent communication.
Conflicting development trends. The first reason derives from the fundamental conflict between the refined development of agents and the abstract demands of users. With the in-depth and specialized evolution of agents towards vertical fields (such as medical diagnostic agents, financial risk control agents, and industrial control agents), their capability boundaries are becoming increasingly refined. However, users’ usage habits exhibit opposite characteristics: they tend to input simple and abstract instructions (such as “plan a cross-border travel”) to trigger the execution of complex tasks. It is hard for a domain-specific agent to finish such abstract instructions independently. Besides, this trend of users’ habits is hard to reverse. This is because people always prefer applications that are easy to operate rather than those that require cumbersome usage steps. The latter is at a disadvantage in the market competition. Usually, for each additional operation step, the user churn rate increases by 10% - 20%. Therefore, agents should not only NOT ask users to make changes, but also cater to such demands of users, which is contrary to the development direction of agents.
Closed ecosystem. The current mainstream multi-agent systems adopt a closed ecosystem design and rely on private interaction mechanisms, forming a rigid technical barrier. This development paradigm severely limits the dynamic collaboration capabilities with external systems. For example, external agents cannot be actively discovered or invoked, making it extremely hard for cross-platform collaboration. Besides, such a closed ecosystem further triggers a decline in systemic effectiveness. On one hand, agents lose scalability due to binding private tools. On the other hand, the obstruction of cross-platform knowledge sharing suppresses the intelligence of agents. As a result, it is necessary to propose agent communication mechanisms that integrate agents from different communities.
IV-B Agent Communication Definition

### v4

@@ II-C Detailed Comparison with the Most Relevant Surveys
Survey [yang2025survey] is one of the surveys with the highest relevance score because it focuses on agent communication and analyzes related protocols. However, it still has significant differences from our survey. First, it lacks some critical protocols like AG-UI, ACP-AgentUnion, ACN, Agent Protocol, API Bridge Agent, and Function Calling. Second, it lacks the analysis of security threats. Third, its defenses prospect is limited. Survey [krishnan2025advancing] focuses on the influences of MCP. As a result, it lacks other protocols, the illustration of agent communication, and security-related discussion. Similarly, the survey [singh2025survey] also focuses on MCP. It lacks illustrations of other protocols and comprehensive security analyses. Survey [ray2025review] comprehensively introduces A2A, lacking discussion about other protocols and security analyses. Survey [ehtesham2025survey] detailed discussed MCP, A2A, ANP, and ACP(-IBM). It also partially analyzed related security risks and defenses. However, there still lacks other protocols, in-depth security analyses, and a systematic illustration of agent communication. Hou et al. [hou2025model] and Zhao et al. [zhao2025mcp] discussed the security risks of MCP. They did not consider other protocols and the high-level overview of agent communication.
III Preliminary: The Evolution of Communication and Agents
In this section, we review the entire lifetime of communication and LLM-driven AI agents.
III-A The Evolution of Communication
As shown in Figure 3, we divide the evolution history of communication technology into four stages: (1) computer communication, (2) mobile communication, (3) IoT communication, and (4) agent communication. These four stages exhibit a close evolutionary relationship characterized by technological inheritance, progressive enhancement of capabilities, and expansion of application scenarios. Each stage lays the foundation for the subsequent one, while integrating the technological advantages of the previous stages.
III-A1 Computer Communication Era
As the starting point of modern communication, this era centered on the interconnection between fixed computer devices. In its initial stage, it relied on dedicated line connections. Later, with the popularization of standard protocols like TCP/IP [fall2012tcp], a global Internet architecture was formed. Its core value lies in building complete standard communication protocols, thus realizing the remote transmission between different types of devices and breaking down the information barriers caused by geographical distances. The main subjects of communication at this stage were hardware devices, only meeting the basic need of data transmission.
III-A2 Mobile Communication Era
With the popularization of mobile communication technologies (from 2G to 6G) [andrews2014will] and intelligent terminals (such as mobile phones), communication scenarios have expanded from fixed computer communication to mobile communication. Compared to the computer communication era, this era has broken through the constraints of time and space, enabling real-time interactions between moving people and devices. The core features are the portability of mobile terminals and the of communication, which has driven the explosive growth of scenarios such as social networking, e-commerce, and mobile office.
III-A3 IoT Communication Era
With the emergence of Internet of Things (IoT) technology [atzori2010internet], the subjects of communication have further expanded to include various IoT devices (such as industrial sensors, cameras, smartwatches, and smart furniture). Communication at this stage is characterized by “thing-to-thing” connectivity. Through customized protocols, real-time perception of the physical world is realized. As a result, the communication has shifted from “human-computer interaction” to environmental perception, providing underlying data support for fields such as industrial automation and smart cities. However, interactions are still mainly focused on data collection and instruction issuance, lacking autonomous decision-making capabilities.
III-A4 Agent Communication Era
Agents’ unprecedented capabilities of reasoning and using tools are expected to make them the foundation of the next communication era. Agent communication has achieved a qualitative leap from device interconnection to collaboration among intelligent entities. Unlike the previous three stages, its communication subjects are LLM-driven agents with capabilities of autonomous perception, memory, reasoning, and decision-making. The communication goal is no longer simple data transmission, but rather revolves around complex tasks involving intention alignment, task decomposition, resource coordination, and result aggregation. It can independently discover collaboration partners, negotiate interaction rules, and dynamically adjust communication strategies. This stage breaks the limitations of devices and scenarios and is expected to become the fundamental support for the future AI ecosystem.
III-B LLM-Driven AI Agents
III-B1 Large Language Model
@@
Large Language Model (LLM) is a new type of artificial intelligence (AI) model trained on large-scale text corpora to understand and generate human language [nvidiallm]. Once it came out, LLMs have demonstrated unprecedented capabilities across a wide range of domains, including but not limited to natural language understanding and generating [yuan2022wordcraft], logic reasoning [qin2023chatgpt, zhong2023can, wei2022chain], code generation [zhang2023planning], and translation [peng2023towards]. These remarkable performances can be attributed to two major factors. One is that LLMs are built upon a powerful architecture known as the Transformer [vaswani2017attention], which effectively models and captures contextual dependencies between tokens and dynamically weighs the importance of different parts of the input. The other key factor, perhaps the most important one, is the massive scales of LLMs that far exceed traditional AI models. When model parameters surpass certain thresholds, LLMs exhibit emergent abilities [wei2022emergent], referring to unexpected capabilities that do not appear in smaller models. As shown in Table II, the parameter scale of an LLM can be hundreds or thousands of times that of traditional AI models.
@@
| FNN | LeNet-5 [lecun1998gradient] | 1998 | 60K |
| RNN | Elman Net [elman1990structure] | 1990 | 100K |
| LSTM | LSTM [hochreiter1997lstm] | 1997 | 1–10M |
| CNN | ResNet-50 [he2016deep] | 2015 | 25M |
| CNN | AlexNet [krizhevsky2012imagenet] | 2012 | 60M |
| CNN | VGG-16 [simonyan2014very] | 2014 | 138M |
| GAN | DCGAN [radford2015unsupervised] | 2016 | 4M |
| GNN | GCN [kipf2016semi] | 2017 | 23K |
| Autoencoders | DAE [vincent2008extracting] | 2008 | 100K |
| Autoencoders | VAE [kingma2013auto] | 2013 | 1M |
| Transformer | GPT-3 [brown2020language] | 2020 | 175B |
| Transformer | PaLM [chowdhery2022palm] | 2022 | 540B |
| Transformer | GPT-4 [achiam2023gpt] | 2023 | 1T |
| Transformer | DeepSeek-V3 [liu2024deepseek] | 2024 | 671B |
| Transformer | DeepSeek-R1 [guo2025deepseek] | 2025 | 671B |
III-B2 LLM-Driven AI Agents
@@
Perception module. To automatically finish a specified task, agents need the ability to perceive the real-world environment. For example, the autonomous driving agent needs to sense road conditions in real time so as to take actions such as avoiding, driving, or braking [mao2023language, ma2024dolphins]. The type of perception ability depends on the domain for which the agent is designed. For instance, an autonomous driving agent needs the ability of visual or radar perception [sun2020mimo, yao2025exploring], while a code-generating agent may not require such functions [huang2023agentcoder, ishibashi2024self].
Memory module. The processing of real-world tasks also requires a strong memory. Agents need to have long-term memory to store complex instructions, knowledge, environment interaction history, or other data that may be required in future steps [hatalis2023memory, maharana2024evaluating, yu2024finmem]. This usually requires external storage resources to assist the brain, such as databases or memory sharing [gao2024memory, gao2025efficient]. In contrast, LLMs do not have such excellent memory ability. Their memory is short-term, which only lasts for rounds of conversations [wang2025recursively, zhong2024memorybank].
Reasoning and planning module. LLM acts as the brain of agents due to its excellent capability of reasoning and planning. It intercepts the instructions from users and automatically decomposes the received task into multiple feasible steps [wang2025tdag, jeyakumar2024advancing, khot2022decomposed, singh2025adaptbot]. Then, it selects the best plan from different candidates [kannan2024smart, zhao2023large, hu2023tree]. Besides, it also revises strategies based on environmental feedback, mitigating errors like code bugs or logical inconsistencies [zhou2024isr, valmeekam2023can, renze2024self, wuproposal, 11209370]. For example, when the autonomous driving module finds that the barrier is closer, it will change the plan to slow down or detour.
Tool module. The tool module is responsible for deeply integrating external resources with the cognitive capabilities of the agent, enabling it to perform complex operations beyond the native capabilities of LLM [yuan2024easytool, wu2024avatar, li2024review, lu2024toolsandbox]. For example, through predefined functional interfaces and protocols, a math agent is able to invoke the external computation libraries and symbolic solvers to help it solve mathematical problems [gou2023tora].
Action module. The action module is the core hub for interaction with the environment. It is responsible for converting the decisions made by LLMs into executable physical or digital operations and obtaining feedback [wang2024executable, zhang2024towards]. This module ensures the executability of instructions through structured output control. For example, it immediately stops generating when LLMs generate a complete action description to avoid redundant output interfering with subsequent parsing.
@@
III-B3 Comparison Between Agents and LLMs
@@
High Autonomy. LLMs can only passively react to the user prompts and then generate responses. They are unable to plan or execute tasks independently. Besides, the response quality highly relies on the prompt skill [white2023prompt, giray2023prompt, mesko2023prompt, liu2023pre, ekin2023prompt, chen2023unleashing, zhou2022learning], which seriously affects the user experience. In contrast, agents possess independent capabilities for task decomposition, strategy adjustment, and external tool invocation, which breaks through the passive mode of LLMS and is highly autonomous.
Flexible Multimodal interaction. LLMs have limited capability of handling multimodal inputs, such as text and pictures [sarch2024vlm, zhang2024vision, zhang2021vinvl, zhou2022learning, laurenccon2024matters, zhu2023minigpt]. Besides, their outputs are also mainly single-modal (e.g., text-only or picture-only), lacking the ability to actively invoke tools to perform physical actions or generate multimodal content. In contrast, agents overcome these drawbacks by deploying multimodal perception frameworks and tool invocation interfaces. They can realize interactions with complex environments, including vision, text, voice, and other physical elements.
Abundant Tool invocation. LLMs usually passively invoke a single tool (such as Function Calling [openai_functioncalling]) through predefined API interfaces and can only perform fixed operations as instructed (e.g., calling the weather API to answer queries [zhang2023ecoassistant]). In contrast, agents have the capability of active decision-making. They can independently select, combine, and dynamically adjust multiple tools, such as connecting crawlers, databases, and visualization tools, to generate responses [he2025pasa].
Better Hallucination inhibition. LLMs suffer from a serious problem called hallucination, which refers to that LLMs are likely to generate non-existent knowledge [zhang2023siren, gunjal2024detecting, tonmoy2024comprehensive, huang2025survey, xu2024hallucination, li2023evaluating]. LLMs mainly rely on the knowledge internalization of training data, making them prone to hallucinations when facing uncovered domains or outdated information. In contrast, agents are able to reduce the error rate by integrating multiple techniques such as Retrieval Augmentation Generation (RAG) [gao2023retrieval, lewis2020retrieval, zhao2024retrieval] or other methods, which can align the action of agents [tennant2024moral, gao2024aligning].
Dynamic adaptability. Essentially, LLMs are static models whose knowledge is fixed at the training phase. Although techniques such as fine-tuning [zhang2024scaling, lin2024data, hu2023llm] or model editing [yao2023editing, wang2024knowledge, zhang2024comprehensive, wang2023easyedit, li2023unveiling] reduce the training cost significantly, LLMs still cannot adapt to real-time events well. In contrast, agents are equipped with techniques like online web search, database query, or real-time sensors, which enable them to dynamically adapt to the changes of real-time environments and information.
@@
Worse Security. Agents have WORSE security than LLMs, which is a major weakness of agents. This is because LLMs are only capable of outputting text. Even if the outputs contain illegal or discriminatory content, their influence on the real world is limited. In contrast, since agents are endowed with the ability to invoke tools, they can cause substantial damages to the real world, including but not limited to maliciously/wrongly operating machines, poisoning databases, and paralyzing the system. For example, attackers can induce agents to visit malicious websites [kong2025web], causing a wide range of subsequent threats. As a result, it is necessary to concentrate more on the security of agents.
III-B4 Agent Applications
@@
Scientific Research. Agents are increasingly embedded into the research workflow, enhancing ideation, automation, and discovery. Their contributions span multiple disciplines, such as mathematics [lei2024macm, deng2024flowdpo, xie2024mathlearner, wang2025malot], chemistry [ruan2024automatic, darvish2024organa, chiang2024llamp, boiko2023autonomous], biological sciences [xin2024bia, xiao2024cellagent, liu2024tais], and materials Science [kumbhar2025hypothesis, uddin2024ai, kulik2024future]
@@
Technical and Engineering Systems. Agents play a growing role in engineering domains, improving automation, systems, and software intelligence. For example, agents are widely used in software engineering, assisting in code generation, bug localization, verification, and system configuration [wang2025training, chen2025locagent, jain2025r2e, hu2025llm, gatelens2025, lu2025uxagent]. Besides, agents are also popular in game development and simulation [npc_behavior_unity_ml_agents, li2024multiagent]. Embodied intelligence is also another hot topic [chen2023scalable, zhao2023roco].
@@
Social Governance and Public Services Agents are increasingly deployed in sectors focused on public service and human welfare. For example, agents are now widely used in the legal field to help draft contracts, review legal documents, check compliance rules, and analyze cases [liu2024efficient, li2024ai, weng2024llms, matthews2024generative]. Besides, other fields, such as financial services [li2024econagent, yang2025twinmarket, okpala2025agentic, yu2024fincon, han2024enhancing, zhang2025finsphere, fatouros2025marketsenseai], education [wang2024ai_education_review, swargiary2024impact, maity2024generative, dickey2024gaide, dumbuya2024personalized, moller2024revolutionising], and healthcare [chen2024cod, wang2025medagent, feng2025m3builder, rose2025meddxagent, averly2025liddia, li2024drugagent, lin2024patentagent, zhang2025psyche, zhang2024psydraw], are also actively integrating agents into their respective practices.
@@
Overall, it can be seen that agents are being widely applied in all walks of life, greatly promoting the development of productivity. More importantly, the application of agents is still in its infancy and has an even greater space for development in the future. It is estimated that the agent market will grow at a rate of 40% annually and is expected to exceed 216.8 billion dollars by 2035 [agengmarket2].
III-C Takeaways
@@
Agents show multiple advantages over LLMs on multiple metrics, such as richer perception ability, stronger learning ability, and higher adaptivity. Now, to improve the service quality, agents are evolving towards refinement to obtain professional skill in a small domain, no longer pursuing the comprehensive capabilities like LLM. LLMs are more like an intermediate transitional form of the future intelligence, while agents are the next stage of development direction of artificial intelligence. It can be foreseen that they will ultimately become indispensable components of future production ecosystems and daily life. However, agents show worse security than LLMs due to their capability to execute tools. As a result, studying the security of agent communication is significant to the AI ecosystem.
IV Agent Communication: Definition and Classification
IV-A Agent Communication Definition